OpenStudy (anonymous):

If the arc on a particular circle has a length of 12 units, and the circumference of the circle is 36, which is the degree measure of the arc?

OpenStudy (calculusfunctions):

The length L of an arc of a circle with radius r and angle θ in degrees is given by\[L =\frac{ \theta }{ 180° }\pi r\]Remember that an arc is part of the circumference of a circle. I'm sure you know that the circumference C of a circle with radius r is given by\[C =2\pi r\]. Now @claudia.valdez95 you should be able to solve the problem on your own. Note: @solarmist7 we don't give out answers. We teach and guide the person to arrive at the correct conclusion themself. Please erase your answer so you don't get reported.